Please calculate the unemployment rate

given the following information:

Please round to two decimal places in your answer. Do not include the % sign

Total Population: 800,270,378

Homemakers: 29,874,567

Retired:112,687,155

Labor Force: 450,314,672

Under the Age of 16: 68,322,654

Discouraged Workers: 98,678,228

Employed: 425,482,365

Underemployed: 14,977,312
Answer #1

Unemployment rate = Number of people unemployed/Labor force*100

Unemployed = Labor force - employed = 450314672-425482365 = 24832307

Unemployment rate = 24832307/450314672*100 = 5.51